OpenStudy (anonymous):

how do i find the slope line 4y - 5 = 6x

OpenStudy (amistre64):

you take the 6 and divide it by the 4

OpenStudy (anonymous):

4y = 6x + 5 divide both sides by 4 y = (6x+5)/4

OpenStudy (anonymous):

what happens to the y and the x?? which variable is the slope

OpenStudy (anonymous):

use y=mx+b so, y will always be dependent mx... m is a number and x is your slope,

OpenStudy (amistre64):

no variable is the slop

OpenStudy (amistre64):

If you want to "find" the slope; get 2 known points from the equation and determine the distances between them

OpenStudy (amistre64):

the easiest points to find or the y intercept and the x intercept

OpenStudy (amistre64):

4y - 5 = 6x when y = 0: (x,0) x = ? x = -5/6 (-5/6, 0) -------------------------- 4y - 5 = 6x when x = 0; (0,y) y=?? y = 5/4 (0,5/4) find our how far and in what direct the "y" value changes between these points from left to right...... and divide that by the distance and direction between the x points from left to right

OpenStudy (amistre64):

the change in y slope = -------------- the change in x

OpenStudy (amistre64):

y = 0 - (5/4) = -5/4 ------------ --- x = -5/6 - (0) = -5/6 slope = -5/4 1/4 --- = ---- = (1/4)(6) = 6/4 = 3/2 -5/6 1/6

OpenStudy (anonymous):

that was alot to take in, but im getting it a little bit now, thankx 2 the both of you guys for replying thank you.

OpenStudy (amistre64):

keep at it, itll click soon enough :)

OpenStudy (anonymous):

thankx, and i will =)

OpenStudy (anonymous):

hey whats up :P

OpenStudy (anonymous):

hey whats up =P

OpenStudy (anonymous):

1+2=3sum

OpenStudy (anonymous):

8==D

OpenStudy (anonymous):

Perhaps to make it easier, the slope is attached to the x. If the numbers are 2/3, the slope is 2/3. If the number attached to the x is whole, like 3x, then it is understood that as a fraction it would be 3/1. Your slope then would be 3/1, read as RISE over RUN, that would be 3 UP and 1 to the right, since both are positive. keep in mind you need to get everything on one side before you can determine the slope. So, in THIS case: you have to first +5 on both sides (zeroing it out on the left). Now it should read 4y=6x+5. Now divide the 4 from both sides. Now it should look like y=6/4x +5/4. Reduce the fraction on the x to 3/2. Rise over run (slope) is 3/2...up 3 over to right 2. What happens to the x and y? They don't disappear...you still have them.

OpenStudy (amistre64):

there are plenty of shorcuts; but the only way to "find" the slope is thru the drudgery of finding points and calculating distances and directions ;) After a few hundred of them you begin to notice that they are follwing a pattern that you will then be able to recognize and skip all the "finding" stuff :)